# Web vs. LLMs: An Empirical Study of Learning Behaviors of CS2 Students

### v1

@@ Abstract.
Empirical Studies, Information Seeking, Large Language Models, CS Education, Web Search
@@ 1. Introduction
The advancement of generative artificial intelligence (AI) has introduced new tools and methods to support CS education. Large Language Models (LLMs) such as ChatGPT have gained popularity for their ability to assist learners through interactive and personalized guidance [2]. These systems can provide high-quality explanations [18, 9], generate examples, and answer questions in real-time, making them a promising resource for learning programming [14].
However, the path to their adoption for self-directed programming learning is replete with challenges: these tools sometimes generate inaccurate or misleading information, lack the ability to understand nuanced learner needs, and potentially foster overreliance on AI instead of critical thinking and independent problem solving  [29]. Recent studies also highlight the importance of learners developing metacognitive abilities – the capacity to recognize what they know, identify gaps in their understanding, and actively address those gaps – to guide their own learning when using LLMs, but these skills might be challenging to grasp for novice learners of programming concepts. [25, 2].
@@
Thus, the research on whether or not to adopt LLMs in learning contexts is divided: some studies such as [4, 22, 32, 2, 14, 11] indicate their promise while others caution against their use [26, 25]. There is also a paucity of empirical studies comparing learning with LLMs against traditional web-based learning (e.g., using video tutorials, documentation, and websites such as Stack Overflow) – both in terms of quantitative outcomes and qualitative behaviors. As a result, making informed decisions about learning choices becomes hard for both students and educators alike.
@@
In this paper, we begin to address this gap by comparing the effectiveness of learning programming concepts with AI vs. traditional web (NoAI) resources. Specifically, we conducted a within-subject lab study at IIT Kanpur, a large engineering university in India, involving 32 participants (CS2 students) learning new topics in programming concepts with LLMs and traditional web resources. We investigated the differences in the strategies learners adopt in both these information-seeking setups, along with their underlying reasons, and the effectiveness of their learning in terms of assessments. Our results suggest that the participants performed better on the post-study evaluations in the NoAI condition. We also observed a statistically significant difference in between the types of prompts in AI and queries under NoAI conditions.
@@
Our key contributions are: 1) a comparison of students’ information-seeking behavior in learning tasks for two programming concepts via a lab study, and 2) an assessment and comparison of students’ understanding of the concepts through quizzes and debugging tasks when learning through AI and NoAI.
@@ 2. Related Work
Prather et al. [22] reviewed the usage of LLMs in computing education, noting their potential to improve learning experiences while cautioning against overreliance [26]. Early experiments with AI-based feedback mechanisms have shown promise in delivering helpful and high-quality feedback[4, 19, 15, 28], although challenges remain to balance automated assistance with learning objectives. Kazemitabaar et al. [14] categorized novice programming strategies into distinct patterns, finding that approaches combining manual coding with AI assistance yielded better results. Amoozadeh et al. [3] report on trust dynamics between students and generative AI, revealing how perceptions influence adoption and outcomes. Overall, this body of literature calls for the deliberate integration of AI in education to facilitate self-regulation and programming skills development.
@@
ChatGPT’s conversational interface has profoundly impacted programming education, as evidenced by the findings of Sun et al. [32]. Amoozadeh et al.[2] discuss how this interface enables personalized learning experiences and promotes student agency. Studies by Kazemitabaar et al.[14] and Guo et al.[11] highlight ChatGPT’s effectiveness in debugging and concept clarification. Although Sun et al. [33] demonstrated that structured prompts improve learning outcomes, Nguyen et al. [20] identified specific barriers in prompt formulation and evaluation of solutions. This duality resonates with Prather et al.’s [23] findings that AI tools accelerate learning for some students while potentially hindering others. While such studies (as listed by Prather et al. [24]) have evaluated the use of LLMs as learning tools, they do not evaluate how learning with LLMs compares to using traditional self-study methods. A notable exception is Qureshi et al.’s study [27] comparing the results of a programming exercise between two groups of students: one that had access to ChatGPT and another that only had access to offline course material. Our study complements Qureshi et al.’s study by comparing ChatGPT with traditional self-study web resources for learning programming concepts.
Online Information Seeking in Programming.
Research on information seeking in the web has identified fundamental patterns in how developers learn programming online. Brandt et al.[7] identified the pattern of “opportunistic programming” demonstrating how developers adapt online examples to their needs. Sellen et al.[30] characterized online information seeking and gathering behavior, providing a framework to understand search strategies. Ko et al.[16] extended this understanding by cataloging information needs during software development. Upadhyay et al.[34] contributed insights into how interface modality affects the information seeking behavior. These studies help us understand how programmers use and navigate between information sources on the web. Our work corroborates and extends this research by examining how online information-seeking behavior differs when using AI tools, providing insights into how these different approaches support programming learning in different ways.
@@ 3. Research Questions
Our specific research questions are:
@@
    Are there any differences, in terms of learning outcomes, between AI and NoAI treatments?
@@
    Are there differences in information seeking strategies in AI and NoAI treatments?
@@ 4. Methodology
To answer the above questions, we conducted a user study with 32 participants, comparing their learning of programming concepts with and without AI. The study was approved by the Institutional Ethics Committee at IIT Kanpur, and all studies were conducted by the first author in January-February, 2024.
@@ 4.1. Study Design
Our study used an in-vitro, within-subject design. Each participant learned two different programming concepts, namely currying  [12] and immediately invoked function expressions (IIFE)  [10] in Javascript. They were instructed to learn one of these tasks using only ChatGPT (GPT 3.5 Turbo) and the other using traditional online resources (e.g web search, videos) starting with a Google search page, without the use of any Generative AI tools. We call these two conditions AI and NoAI respectively. Participants in both conditions had access to the Visual Studio Code (VSCode) IDE to run code if they decided. The task assignment to conditions and task order were balanced.
4.2. Participants and recruitment
In December 2023, we sent a recruiting email to the students’ mailing list at IIT Kanpur inviting students to participate in a 1-hour study. The email included a sign-up form which contained questions about respondents’ programming experience, programming courses taken, programming languages known, and generative AI usage. We received 71 responses before closing replies to the form. We then invited 32 respondents: 31 were undergraduate students who had taken CS1 and were eligible for CS2, and 1 was a non-CS graduate student (details of participants’ majors are in Table 2) All participants were familiar with programming, and had prior experience using Generative AI (GenAI) tools, but did not know Javascript (Table 1).
@@ 4.2. Participants and recruitment
Table 1. Participants’ reported prior usage of GenAI tools
Computer Science Engineering Sciences and Others
6 20 6
Table 2. Participants’ major field
@@ 4.3. Study Protocol
Each study session started with a briefing and informed consent. Participants were then given 15 minutes to to learn the concept (with or without AI) followed by an assessment. They then moved on to the learning and assessment for the second task/condition. Participants were asked to think aloud during the learning task. For the assessment, participants took a conceptual-understanding quiz followed by a 10-minute debugging task. Participants were informed about assessments during the recruitment and pre-study briefing. At the end of the study session, participants were asked questions about their fixes in the debugging task to determine if their task completion (if any) was accompanied by an understanding of the program, the bug, and its fix, thus indicating learning. Participants received a cash compensation of INR 150 for their participation.
@@ 4.4. Tasks and Assessments
4.4.1. Learning tasks.
Participants learnt two programming concepts associated with functional programming . All the pre-requisites for understanding these concepts (such as functions and variable scopes) were part of a mandatory CS1 course that all participants had taken. For each task, participants were provided with the following learning objectives to guide their learning: ”You should be able to answer the following questions through your learning: What is ¡concept¿ in programming? Why is ¡concept¿ used in programming? How is ¡concept¿ implemented in programming languages?” . This was to prevent participants from being derailed by the historical, mathematical or programming language implementation aspects of the concepts, as was the case in pilot studies. Since Javascript was the language used in our study, we ”recommended” that participants ”use JavaScript as the medium to understand the concepts”.
@@
We evaluated participants’ learning of each task using a two-part assessment. The first part assessed conceptual understanding via a 5-question quiz, comprising of 4 true/false questions related to the syntax and purpose of the concept and 1 multiple choice question about the benefits of applying the concept. All questions also had a ”don’t know” option and participants were instructed before the quiz to avoid guessing and choose this option if they did not know the answer to a particular question. The second part of the assessment involved participants debugging and fixing a semantic error in a Javascript program to work according to the stated specification. The code only contained rudimentary Javascript structures and did not contain Javascript-specific idioms. A patch could contain up to five lines of code. We set a 10-minute time limit for this task based on pilot studies.
@@ 4.5. Data Collection and Analysis
4.5.1. Participants’ artifacts
We gathered participants’ solutions to the two assessments (i.e., quiz and debugging task) and scored them for further analysis.
For the quiz, each true/false question was awarded +1 or -1 depending on the correct / incorrect answer, while the multiple choice question (which contained multiple correct answers) was scored as +1, -1 based on all correct and incorrect answers respectively. A ”don’t know” response for any question received a score of 0. Thus, the possible quiz scores ranged from -8 to +8, with incorrect understanding carrying a greater penalty than ”I don’t know”.
For the debugging task, we assigned three possible scores: 1) score=0, if the final code failed to produce desired output, 2) score=1 if the final code produced the desired output and the participant was able to correctly describe their bug fix, and 3) score=0.5, if the code produced the desired output but the participant was unable to explain their solution. We refer to these outcomes as Incorrect, Correct and Correct (LoU), where LoU stands for lack of understanding.
Category Code Definition Example Programming Coding Participant uses the VSCode IDE while learning Participant writes and/or executes code on VSCode Video Locating Participant looks for a video to watch Searching ”javascript currying tutorial” on YouTube Learning Watching Participant learns by watching a video Watching a tutorial explaining IIFE implementation Learning Locating Searching for information to learn from Scrolling through search results for a search query; Writing a prompt to ChatGPT Learning: Theory Learning through abstract theory Reading the definition and advantages of currying Learning: Example Learning through information related to examples Reading the description of an example of an IIFE
@@ 4.5.1. Participants’ artifacts
Code Definition Example Query Origin Copied Queries directly pasted from external sources without modification Copying ”What is currying in programming?” from the learning objectives Edited Queries edited or written completely by participants Modifying the original query to ”What is currying in javascript?” Follow-Up Base Queries starting a new line of thought independent of previous queries or learned information What is currying in javascript? Derived Queries associated with a previous query or learned information that continue exploring an existing line of thought, such as follow-ups, modifications or clarifications Explain this example in more detail Query Phrasing Evidence-Based Queries that ask about the features/description/definition of a concept/idea/object/event What is currying in programming? Reason You want to find out reasons of/for something. Why are IIFEs used? Comparison Queries that ask to compare/contrast two or more things, understand their differences/similarities. How are curried functions better than normal functions? Instruction Queries to understand the procedure/method of doing/achieving something. How do I write an IIFE in javascript? Debate Queries that debate a hypothetical question (is someone right or wrong, is some event perceived positively or negatively?). Do you really think currying can be useful? Experience Queries that ask for advice or recommendations on a particular topic. Do you recommend using currying? Statement Queries not phrased as a question that do not solicit information Currying in javascript Explanatory Queries not phrased as a question that solicit information Please explain how currying helps with function composition
Table 4. Query Classification Codebook
4.5.2. Task videos
We recorded participants’ video, audio and screen actions during the study. We qualitatively analyzed this data in two parts. First, we performed activity analysis to understand the time participants spent on different learning activities (Table 3). For this, we coded the videos in 30-second segments, with one or more codes per segment. Our codes consisted of programming, information seeking, and information learning [31], where learning was divided into theoretical learning and learning by example [7]. We also recorded the number of videos watched and websites visited by each participant. Second, we performed a qualitative analysis of search queries and LLM prompts to understand their nature (Table 4). We collected all queries and prompts along with their timings. We categorized them across three dimensions: follow-up, origin, and phrasing. As web searches often lead to follow-up queries [31], this analysis helps us examine if and how LLM prompts lead to follow-ups. The taxonomy for query phrasing has been adapted from Bolotova et al.’s [6] taxonomy for questions. We extended this taxonomy to queries not expressed as questions by adding two categories - Statement and Explanatory.
4.5.3. Reliability.
Two authors coded 4 out of 32 (¿10%) videos independently to design an initial codebook. The interrater reliability was 87% on the Jaccard similarity index for activity analysis (Table 3) and was 89%, 83% and 81% on Cohen’s Kappa [8] for the dimensions of origin, follow-up, and phrasing, respectively, in the query/prompt analysis (Table  4). We used the Jaccard index for activity analysis as the codes were not mutually exclusive and multiple code assignments per segment were allowed. Following this, the remaining videos were divided and coded by one of the two coders based on the agreed upon codebook.
5. Results
5.1. RQ1: Post-study assessment results
5.1.1. Debugging Task Performance.
Figure 2 gives an overview of the results of the debugging task across tasks and treatments. The paired t-test did not find significant differences across the dimensions of concept learned (p-value = .08), or learning treatment (p-value = .57). We also did not find any significant differences on applying the Mann-Whitney U-test across the dimensions of learning treatment when considering the debugging task performance for currying (p-value = .53) and IIFE (p-value = .88).
Figure 1. Debugging Task Performance
Figure 2. Theory Quiz Scores
5.1.2. Quiz Performance
Figure 2 presents the distribution of quiz scores for each concept and learning treatment, scored on a scale of [-8, +8]. We found that the quiz scores for currying (M = 0.66, SD = 2.52) were significantly lower (p = 0.0002) than IIFE (M = 2.69, SD = 2.10). The effect size, measured by Cohen’s d, was d = -0.86, indicating a large effect. It also took participants more time to complete the currying quiz compared to IIFE (163.1 vs 97.21 seconds, on average, paired t-test p-value ¡¡ .05). Across all participants and questions in the quiz, participants chose the ”don’t know” option 31 times (15.9% of all options chosen) in the quiz for currying, and 29 times (13.5% of all options chosen) in the quiz for IIFE. This may suggest that currying is a more difficult topic for participants to learn in the study condition than IIFE in general, or that the post-study quiz in currying task was harder than in IIFE. As a result, we compare the NoAI and AI conditions for each task.
AI vs. NoAI in Currying
Our sample size was small (N = 16) and the currying quiz scores for the AI treatment were non-normal [Shapiro-Wilk (W = 0.84, p = 0.01)]. Therefore, we conducted a Mann-Whitney U-test and found statistical differences between the mean quiz scores in the AI and NoAI treatments (-0.69 vs. 2.0, p=.005). The effect size was large (r = 0.50).
AI vs. NoAI in IIFE
In contrast, we did not find any significant difference in the IIFE quiz scores between the AI and NoAI conditions (p = 0.34). Here, we used an independent sample t-test since Shapiro-Wilk tests suggested normal distributions of IIFE quiz scores in AI (W = 0.92, p =.20) and NoAI (W = 0.93, p = .28) treatments.

### v3

# To Google or To ChatGPT? A Comparison of CS2 Students’ Information Gathering Approaches and Outcomes

@@ Abstract.
copyright: none
@@ 1. Introduction
With the widespread adoption of digital education [45], understanding the tools and processes behind self-directed (SDL) and self-regulated learning (SRL) has become increasingly important. Online educational resources tend to be the primary tools for self-directed learners [44], especially in the case of learning programming [23]. The advancement of generative artificial intelligence (AI) has introduced a variety of online tools and methods to support SDL [47]. In particular, Large Language Models (LLMs) such as ChatGPT have gained popularity for their ability to assist learners through interactive and personalized guidance [2]. Unlike searching through the web, which often requires complex navigation between multiple sources [36], these models respond directly to user queries in real-time. They can generate examples as well as high-quality explanations [20, 9], making them a promising resource for learning programming [15]. However, the path to their adoption for self-directed programming learning is replete with challenges: these tools sometimes generate inaccurate or misleading information, lack the ability to understand nuanced learner needs, and potentially foster over-reliance on AI instead of critical thinking and independent problem solving  [35]. Thus, research on whether or not to adopt LLMs in learning contexts is divided: some studies such as [4, 29, 41, 2, 15, 12] indicate their promise, while others caution against their use [32, 31].
@@
As the number of online resources for learning programming increases, so does the complexity of information problems for self-directed learners [25]. However, there is a paucity of empirical studies comparing information gathering approaches with LLMs to traditional web-based learning (e.g., using video tutorials, documentation, and websites such as Stack Overflow) – both in terms of quantitative outcomes and qualitative behaviors [38]. As a result, making informed decisions about learning choices becomes difficult for both students and educators alike.
@@
In this paper, we begin to address this gap by comparing how learners gather information about programming concepts with AI vs. traditional web (NoAI) resources such as Google search and YouTube videos. Specifically, we conducted a within-subject lab study at IIT Kanpur, a large engineering university in India, involving 32 participants (CS2 students) learning new topics in programming concepts with LLMs and traditional web resources. We investigated the differences in the strategies learners adopt in both these information-seeking setups, along with their underlying reasons, and the effectiveness of these strategies in terms of assessments. Our results suggest that the participants performed better on the post-study evaluations in the NoAI condition. We also observed a statistically significant difference between the types of prompts in AI and queries under NoAI conditions.
@@
Our key contributions are: 1) a comparison of students’ information-seeking behavior in learning tasks for two programming concepts via a lab study, and 2) an assessment and comparison of the effectiveness of students’ information-seeking strategies through quizzes and debugging tasks when using AI and NoAI.
@@ 2. Related Work
A review of research on student interactions with AI tools in programming education reveals both opportunities and challenges. Prather et al. [29, 32] reviewed the use of LLMs in computing education, highlighting their potential to enhance learning while warning against overreliance. AI-based feedback mechanisms show promise in supporting students [4, 21, 16, 34], though balancing automation with pedagogy remains a challenge. Kazemitabaar et al. [15] categorized novice programming strategies and found that hybrid approaches (manual coding + AI assistance) yield better results. Amoozadeh et al. [3] examined trust dynamics, showing how perception influences AI adoption. Researchers agree that integrating AI in a structured and deliberate manner can enhance self-regulation and programming skills for CS students in higher education [29]. This deliberate integration calls for an understanding of how students gather information when learning with LLMs and the scenarios in which their strategies are most effective.
@@
The conversational interface of ChatGPT can enhance programming education by enabling personalized learning, as shown by Sun et al. [41] and Amoozadeh et al. [2]. Studies by Kazemitabaar et al. [15] and Guo et al. [12] highlight its ability to aid debugging and concept clarification. Prior research has indicated that prompt strategies play a key role in determining the effectiveness of ChatGPT. Sun et al. [42] reported that structured prompts improve outcomes, while Nguyen et al. [27] identified challenges in prompt formulation and solution evaluation. Lee et al. [19] found that generative AI shifts effort from searching to verifying, requiring stronger prompt and validation skills. This information-seeking gap is reflected in Prather et al.’s  [30] findings that while AI accelerates learning for some students, it may hinder others. Skripchuk et al. [38] investigated student intentions behind using the web or LLMs to learn programming, highlighting that comparisons of information-seeking strategies between LLM-based and traditional self-study methods remain scarce. A study by Qureshi et al. [33] found students using ChatGPT outperformed those relying solely on offline materials for a programming assessment. Our study extends this line of work by comparing AI-assisted and web-based self-study in programming education.
Online Information-Seeking Strategies in Programming
Prior research on information-seeking behavior has identified key patterns in how developers learn programming online. Sellen et al. [36] classified information gathering tasks as a core aspect of knowledge work for developers, accounting for 35% of web activities. These tasks often require navigating multiple sources to gain a holistic understanding. Ko et al. [17] mapped common developer information needs, while Upadhyay et al. [43] explored how interface modality affects search behavior. Duala-Ekoko and Robillard [10] identified three primary information-seeking strategies in programming learning: web-based (seeking code online), documentation-based (official manuals), and hybrid (combining both). Prior work has also described how developers manage information-seeking tasks through strategies such as opportunistic programming  [7], in which developers adapt online examples to their needs. With AI emerging as an important online resource for programmers [39], there is a need to understand how these strategies compare to those of gathering information with AI tools.
AI and Web-Based Resources in Self-Directed Learning
Self-directed learning (SDL) is an essential part of programming education, fostering problem-solving skills and adaptability  [46]. Computing students engage in SDL through personal projects, social influences, and skill competitiveness  [22]. AI-driven tools are reshaping SDL by providing personalized instruction and automated feedback  [40]  [26]. Younas et al. [48] review AI-based SDL approaches, highlighting key tools such as adaptive tutoring and real-time feedback. Structured AI interactions enhance engagement, as seen in ChatGPT-based learning environments [40]. While prior research indicates the promise of AI for SDL, it is not without its obstacles. Wirth et al. [46] suggest that the effectiveness of AI tools depends on structured pedagogy, ensuring students engage critically rather than relying passively on AI. Studies indicate that students with prior knowledge and digital literacy benefit the most from AI-based SDL, highlighting the need for structured scaffolding [48]. Examining learners’ information-seeking strategies with AI in the context of traditional methods can facilitate this structured adoption of AI by providing insights into how different tools benefits students in a self-directed learning environment.
@@ 3. Research Questions
We aim to address the following research questions:
@@
    What, if any, are the differences in information-seeking strategies between the web and AI treatments?
@@
    How effective are the information-seeking strategies in the web and AI treatments in terms of conceptual understanding and practical application?
@@ 4. Methodology
To answer the above questions, we conducted a user study with 32 participants, comparing their information-seeking strategies when learning programming concepts with and without AI. The study was approved by the Institutional Ethics Committee at IIT Kanpur.
@@ 4.1. Study Design
Our study used an in-vitro, within-subject design. Each participant learned two different programming concepts, namely currying  [13] and immediately invoked function expressions (IIFE)  [11] in JavaScript. They were instructed to learn one of these tasks using only ChatGPT (model version GPT 3.5 Turbo), and the other using traditional online resources (e.g web search, videos) without the use of any generative AI tools starting with a Google search page. We call these two conditions AI and NoAI respectively. Participants in both conditions had access to the Visual Studio Code (VSCode) IDE to run code if they decided. The task assignment to conditions and task order were balanced.
Computer Science Engineering Sciences and Others
6 20 6
Table 1. Participants’ major field (n = 32)
@@
Table 2. Participants’ reported prior usage of generative AI tools (n = 32)
4.2. Participants and recruitment
We sent a recruiting email to the students’ mailing list at IIT Kanpur inviting students to participate in a 1-hour study. The email included a sign-up form which contained questions about respondents’ programming experience, programming courses taken, programming languages known, and generative AI usage. We received 71 responses before closing replies to the form. We then invited 32 CS2-eligible respondents: 31 were undergraduate students who had taken CS1 at IIT Kanpur, and 1 was a non-CS graduate student (details of participants’ majors are in Table 1)
All participants were familiar with programming, and had prior experience using generative AI tools (Table 2), but did not know JavaScript.
[…]
Figure 1. Study Protocol (A/B refer to Currying or IIFE; X/Y refer to AI or NoAI)
@@ 4.3. Study Protocol
Each study session started with a briefing and informed consent. Participants were then given 15 minutes to learn about the concept (with or without AI) followed by an assessment. They then moved on to the learning and assessment for the second task and treatment. Figure 1 shows the procedure of each session. For example, a participant would first learn about currying with AI and then learn about IIFE without AI.
In earlier stages of our study protocol, participants were given over 30 minutes to work on one task. However, we observed in pilot studies that given this much time, participants tended to feel frustrated and fatigued if they were unable to make progress. We thus shifted to a 10 minute limit for learning time before extending it to 15 minutes. In subsequent pilot studies, we observed that 15 minutes was sufficient for participants to gather enough information about the concept to solve our assessments.
Participants were asked to think aloud during the learning session. If they forgot, they were prompted to do so during the session. After the learning task, the researcher conducted a retrospective interview. For the assessment, participants took a conceptual-understanding quiz followed by a 10-minute debugging task. Participants were informed about assessments during the recruitment and pre-study briefing. At the end of the debugging task, participants were asked questions about their fixes in the debugging task to determine if their task completion (if any) was accompanied by an understanding of the program, the bug, and its fix, thus indicating knowledge of the application of the concept. Participants received a cash compensation of INR 150 for their participation.
@@ 4.4. Tasks and Assessments
4.4.1. Tasks.
Participants were tasked with gathering information about two programming concepts associated with functional programming. All the pre-requisites for understanding these concepts (such as functions and variable scopes) were part of a mandatory CS1 course (that used C as the language of instruction) that all participants had taken. For each task, participants were provided with the following learning objectives to guide their learning: ”You should be able to answer the following questions through your learning: What is ¡concept¿ in programming? Why is ¡concept¿ used in programming? How is ¡concept¿ implemented in programming languages?” . This was to prevent participants from being derailed by the historical or mathematical aspects of the concepts, as was the case in pilot studies. Since JavaScript was the language used in our study, we ”recommended” that participants ”use JavaScript as the medium to understand the concepts”, but did not mandate it, as we observed in pilot studies that some students preferred to learn about the concept using programming languages they knew before understanding its implementation in JavaScript. Our pilot studies also confirmed that ChatGPT consistently generated technically correct examples and explanations of both concepts.
@@
We evaluated participants’ learning of each task using a two-part assessment. The first part assessed conceptual understanding via a 5-question quiz, comprising of 4 true/false questions related to the syntax and purpose of the concept and 1 multiple choice question about the usefulness of the concept. All questions also had a ”don’t know” option. Participants were instructed before the quiz to avoid guessing and choose this option if they did not know the answer to a particular question. The second part of the assessment involved participants debugging and fixing a semantic error in a JavaScript program to work according to the stated specification. The code only contained rudimentary JavaScript structures and did not contain JavaScript-specific idioms. A patch could contain up to five lines of code. We set a 10-minute time limit for this task based on pilot studies. The quiz and debugging assessments are described in section A and B of the appendix respectively.
@@ 4.5. Data Collection and Analysis
We collected the following artifacts from participants:
  1. (1)
    Screen actions via video and audio of the study session
  2. (2)
    LLM prompts and web queries written during the learning task and their timestamps
  3. (3)
    Number of unique websites visited and videos watched during the web treatment of the learning task
  4. (4)
    Responses to the quiz and debugging assessment
We scored participants’ responses to the two assessments for further analysis.
For the quiz, each true/false question was awarded +1 or -1 depending on the correct / incorrect answer, while the multiple choice question (which contained multiple correct answers) was scored as +1, -1 based on all correct and incorrect answers respectively. A ”don’t know” response for any question received a score of 0. Thus, the possible quiz scores ranged from -8 to +8, with incorrect understanding carrying a greater penalty than ”don’t know”.
For the debugging task, we assigned three possible scores: 1) score=0, if the final code failed to produce the desired output, 2) score=1 if the final code produced the desired output and the participant was able to correctly describe their bug fix, and 3) score=0.5, if the code produced the desired output but the participant was unable to explain their solution. We refer to these outcomes as Incorrect, Correct and Correct (LoU) respectively, where LoU stands for lack of understanding.
5. Results
5.1. RQ1: What, if any, are the differences in information-seeking strategies between the web and AI treatments?
To answer this question, we investigated the information gathering behaviors of participants based on their screen actions during the learning session as well as their prompting/querying strategies across the two treatments. The following subsections lay out how we qualitatively coded the data and the results of this analysis.
Figure 2. Example application of activity analysis codebook (vertical divisions in a segment imply multiple codes applied)
Category Code Definition Example
Programming Coding Participant uses the VSCode IDE while learning Participant writes and/or executes code on VSCode
Video Locating Participant looks for a video to watch Searching ”javascript currying tutorial” on YouTube
Learning Watching Participant learns by watching a video Watching a tutorial explaining IIFE implementation
Learning Locating Searching for information to learn from Scrolling through search results for a search query; Writing a prompt to ChatGPT
Learning: Theory Learning through abstract theory Reading the definition and advantages of currying
Learning: Example Learning through information related to examples Reading the description of an example of an IIFE
@@ 5. Results
Activity Treatment Time Spent (%)
Minimum Maximum Mean
Locating AI 7% 40% 26%
NoAI 7% 38% 22%
Learning AI 16% 80% 51%
(Example) NoAI 3% 73% 40%
Learning AI 0% 40% 15%
(Theory) NoAI 2% 52% 17%
Using IDE AI 0% 37% 8%
NoAI 0% 39% 10%
Watching Videos NoAI 0% 56% 11%
5.1.1. Activity Analysis
To analyse the time participants spent on different information gathering activities, we created a codebook (Table 3) to describe participants’ screen actions. Our codes consist of programming, locating information sources, and information learning [37], where learning was divided into theoretical learning and learning by example [7]. Actions were coded in 30-second segments, with one or more codes per segment. Figure 2 shows an example of the application of this coding scheme. To validate this codebook, two authors coded 4 out of 32 (¿10%) videos independently. The interrater reliability was 87% on the Jaccard similarity index. We use the Jaccard index as the codes were not mutually exclusive and multiple code assignments per segment were allowed.
The distribution of the time spent by participants on various activities is presented in Table 4. In the NoAI treatment, the data for locating information sources includes the data for locating websites (20.7%) as well as videos (1.4%). As the table shows, in both conditions, participants showed a strong tendency to learn by example. P21 eloquently described why they used examples to understand concepts: “With examples, I can try to implement them and by creating variations I can understand the concept better. They are easier to comprehend… not abstract”. Others cited that examples were helpful to understand the applications and usefulness of a concept: ”It wasn’t very obvious why currying was useful, but through the examples it is very clear” (P25).
Likewise, in the NoAI treatment, 25% of all 32 participants watched videos, finding that videos were more accessible than websites and made concept understanding easier. For example, P22 said: ”I felt like these websites are directly jumping to the output… I am not familiar with the syntax so it would be good if I can see the flow, and a YouTube video can help with that”. P28 mentioned that videos can provide a “demonstration” of concepts and P20 mentioned that videos felt more like “human interactions” than websites.
Nearly two-thirds of participants (23 out of 32) chose to try examples in the VSCode IDE while learning, across the AI and NoAI conditions. However, this amounted to only about 10% of their learning time on average (Table 4). This can be attributed to the fact that many participants only used the IDE to run code from external sources and verify its expected output, and did not tinker with them. 7 participants mentioned that they used the IDE only to verify ChatGPT-generated code and 5 mentioned the same for websites.